A snowmobile has a mass of 2.50 × 102 kg. A constant force is exerted on it for 60.0 s. The
snowmobile’s initial velocity is 6.00 m/s and its final velocity 28.0 m/s.
What is its change in momentum?
What is the magnitude of the force exerted on it?
1
Expert's answer
2012-03-20T11:07:36-0400
F=p/t
F=ma=m(V1-V2)/t=91
p=F/t=1.57